OpenStudy (anonymous):

MEDAL REWARDED Find the exact value by using a half-angle identity. sin 7pi/8

OpenStudy (anonymous):

@John_ES

OpenStudy (anonymous):

Do you want the answer choices?

OpenStudy (john_es):

No needed. :)

OpenStudy (john_es):

First thing we must know is that, \[\sin(7\pi/8)=\sin[(7\pi/4)/2]=-\sqrt{\frac{1}{2}(1-\cos(7\pi/4))}=-\sqrt{\frac{1}{2}(1-1/\sqrt{2})}\] \[\sin(7\pi/8)=-\frac{\sqrt{2-\sqrt{2}}}{2}\]

OpenStudy (john_es):

Sorry, without the minus sign.

OpenStudy (anonymous):

All my answer responses have 1/2 infront?

OpenStudy (john_es):

Yes, it should be, \[\sin(7\pi/8)=\frac{1}{2}\sqrt{2-\sqrt{2}}\]